OpenStudy (anonymous):

A warehouse has 5,000 drums of liquid chemicals. The foreman knows that 50 of these drums contain off-quality chemicals. Assume that the drum is replaced each time before sampling. What is the probability that none of 3 drums he selects at random are off-quality?

OpenStudy (anonymous):

P(off quality) = 50/5000 P(quality) = 4950/5000 = 99/100 P(3 drums, quality) = (99/100)(99/100)(99/100) = .970299